### v2

@@ Appendix C Euler characteristic number and Chern number in 2-dimensional system
Given a complex linear space ${}^{\mathbb{C}}V$ with inner product $(\cdot,\cdot)_{\mathbb{C}}$, we can realify it to get a real linear space ${}^{\mathbb{R}}V$ with corresponding inner product $(\cdot,\cdot)_{\mathbb{R}}$
$(X,Y)_{\mathbb{R}}:=\operatorname{Re}(X,Y)_{\mathbb{C}},~{}\forall X,Y\in V.$
So that any vector $X\in V$ has the same norm in both ${}^{\mathbb{C}}V$ and ${}^{\mathbb{R}}V$, i.e. $||X||_{\mathbb{C}}=||X||_{\mathbb{R}}$, which can be guaranteed by polarization identity
$(X,Y)_{\mathbb{C}}=\frac{||X+Y||_{\mathbb{C}}^{2}-||X-Y||_{\mathbb{C}}^{2}}{4}+i\frac{||X+iY||_{\mathbb{C}}^{2}-||X-iY||_{\mathbb{C}}^{2}}{4},$
$(X,Y)_{\mathbb{R}}=\frac{||X+Y||_{\mathbb{R}}^{2}-||X-Y||_{\mathbb{R}}^{2}}{4}.$
As a result, the realification of ${}^{\mathbb{C}}V$ seems like that we “forget” the multiple of imaginary number in $V$.
After realifing the eigen subspace we get a real linear space $W_{n}$ with corresponding inner space (,)($\cdot$,$\cdot$)( ⋅ , ⋅ )
$W_{n}:={\rm span}\{\ket{n(k)},~{}i\ket{n(k)}\},(\alpha\ket{n(k)},\beta\ket{n(k)})=\operatorname{Re}(\bar{\alpha}\beta)\braket{n(k)}{n(k)}.$ (C.7)
It is interesting to consider a 2-dimensional Riemann manifold whose tanget space at $k$ point is $W_{n}$ with the $k\in\mathbb{T}^{2}$ is the local coordinates.
Definition C.2
Give a Hamiltonian $\hat{H}(k)$ in 2-dimensional system, consider one of its energy band $E_{n}(k)$
$\hat{H}(k)\ket{n(k)}=E_{n}(k)\ket{n(k)},~{}k\in\mathbb{T}^{2}.$
Then we can define a close surface with orthogonormal frame $e_{1}=\ket{n(k)},e_{2}=i\ket{n(k)}$, and corresponding measure 1-form is $\theta^{1}=\operatorname{Re}(\bra{n(k)}\cdot),\theta^{1}=\operatorname{Re}(-i\bra{n(k)}\cdot)$, we denote $\theta^{1}=\bra{n(k)}$ | and $\theta^{2}=-i\bra{n(k)}$ | in short, so that $\braket{\theta^{i},e_{j}}={\delta_{j}}^{i}$.
Then the close surface $M_{n}$ is defined as the Riemann manifold with following metric $g$ with $k\in\mathbb{T}^{2}$ is the local coordinates
$g=\theta^{1}\otimes\theta^{1}+\theta^{2}\otimes\theta^{2}.$
After getting the close surface $(M_{n},g)$, it is natural to investigate its connection, curvature and topology, it can be easily checked the Levi-Civita connection 1-form ${\omega_{1}}^{2}=-{\omega_{2}}^{1}=i\bra{n(k)}d\ket{n(k)}$, which satisf Cartan’s structure equation $d\theta^{i}=\theta^{j}\wedge{\omega_{j}}^{i}$. And there is no doubt that the curvature 2-form ${\Omega_{2}}^{1}=-d\bra{n(k)}\wedge d\ket{n(k)}$ (Notice that the inner product needs to be taken the real part). It is to say that the Levi-Civita connection and corresponding curvature form for $M_{n}$ are Berry connection and Berry curvature respectively, which truly have geometric meanings. Using Gauss-Bonnet’s formula we can calculate the Euler characteristic number $\chi$ for $M_{n}$
$\chi=\frac{1}{2\pi}\int_{M_{n}}{\Omega_{2}}^{1}=-\int_{\mathbb{T}^{2}}\frac{d\bra{n(k)}\wedge d\ket{n(k)}}{2\pi}.$ (C.8)
The Euler characteristic number $\chi$ for $M_{n}$ is the Chern number $C(E_{n})$ for energy band $E_{n}$ exactly. If there is any isolated point $k_{0}\in\mathbb{T}^{2}$ where $E_{n}(k_{0})$ degenerate, it will be difficult to choose corresponding eigen vector |$\ket{n(k_{0})}$. It might be natural to choose the |$\ket{n(k_{0})}$ as the limit of |$\ket{n(k)},k\to k_{0}$, which aims to keep the continuity of the eigen vector, but the limit may not be exist, so we claim if there is any zero point for |$\ket{n(k)}$, the energy bands degenerate at some point.
From Poincare-Hopf’s theorem we know that for a tangent vector field on $M_{n}$, the sum of the zero point indices equals to the Euler characteristic number $\chi$ for $M_{n}$, i.e. there must be degenerate point on the energy band when the Chern number for the band is 00.
Chern number, or Euler characteristic number for 2-dimensional system, is topological invariant, which means it will not change under topological transformation. In the word of physics, the Chern number cannot be changed by perturbation. As it has been mentioned, non-zero Chern number means there is degenerate point for the energy band, which will not be gapped by small perturbation.

### v3

@@ Appendix C Euler characteristic and Chern number
Let $(^{\mathbb{C}}V,(\cdot,\cdot)_{\mathbb{C}})$ be a complex vector space. Its realification $^{\mathbb{R}}V$ is the real vector space equipped with the inner product:
$(X,Y)_{\mathbb{R}}:=\operatorname{Re}(X,Y)_{\mathbb{C}}.$ (C.11)
This preserves norms through the polarization identities:
[…]
The realification process effectively ”forgets” the complex structure while maintaining the metric properties.
Through realification, we obtain a real vector bundle $W_{n}$ with fibers:
[…]
equipped with the real inner product:
[…]
This construction naturally leads to a Riemannian interpretation:
Definition C.2 (Quantum State Manifold)
[…]
we define a closed surface $M_{n}$ with:
[…]
  • Riemannian metric: $g=\theta^{1}\otimes\theta^{1}+\theta^{2}\otimes.\theta^{2}$
The geometry of $M_{n}$ reveals deep connections:
  • The Levi-Civita connection 1-form is:
[…]
    satisfying Cartan’s structure equations $d\theta^{i}=\theta^{j}\wedge\omega_{j}{}^{i}$.
  • The curvature 2-form becomes:
[…]
    where the inner product is taken in the realified space.
Applying the Gauss-Bonnet theorem yields:
[…]
establishing the exact equality between the Euler characteristic and Chern number.
Key observations about degeneracies:
[…]
  • The Poincaré-Hopf theorem requires that when $\chi(M_{n})\neq 0$, the vector field must have zeros corresponding to band degeneracies.
  • These degeneracies are topologically protected - they cannot be removed by small perturbations.
The Chern number’s topological nature implies:
  • Robustness against continuous deformations of the Hamiltonian.
  • Quantized Hall conductivity in quantum Hall systems.
  • Protection of edge states in topological insulators.
Appendix D Concise explanation for effectiveness of topological Hamiltonian
The topological Hamiltonian is defined in terms of the zero-frequency Green’s function. It is therefore pertinent to inquire whether potential poles in the complex frequency domain impact the validity of the topological Hamiltonian method. In this appendix, we will clarify why the topological Hamiltonian approach remains valid without requiring that the Green’s function has no singularities or poles in the imaginary frequency space, and justify our use of real-frequency Green’s functions for holographic semimetals.
As derived in the original work proposing the topological Hamiltonian approach [26], the topological invariant in terms of Green’s functions for interacting systems could be calculated from
$N_{2}=\frac{1}{24\pi^{2}}\int dk_{0}d^{2}k\text{Tr}[\epsilon^{\mu\nu\rho}G\partial_{\mu}G^{-1}G\partial_{\nu}G^{-1}G\partial_{\rho}G^{-1}],$ (D.1)
where $G(w,k)$ is the Green’s functions. In the original work, the authors have used imaginary frequency Green’s functions, which are the Matsubara Green’s functions, while emphasizing that their final conclusions are also applicable using real frequency. The reason that they used imaginary frequency Green’s functions is that this could be used at finite temperature and the imaginary frequency Green’s functions are more well behaved in their case. At the same time calculations on real frequency Green’s functions for interacting fermionic systems are much more complicated than imaginary frequency ones in condensed matter physics.
The topological Hamiltonian approach relies on a smooth deformation of the Green’s function
$G(iw,k,\lambda)=(1-\lambda)G(iw,k)+\lambda[i\omega+G^{-1}(0,k)]^{-1},$ (D.2)
where $\lambda\in[0,1]$. They have shown that $G(iw,k,\lambda)$ is a smooth function so tuning from $\lambda=0$ to $\lambda=1$ does not modify the topological properties. Therefore, using the effective Green’s function $i\omega+G^{-1}(0,k)]$ (corresponding to $\lambda=1$) would give the original and correct topological invariant (corresponding to $\lambda=0$). Then substituting the $\lambda=1$ Green’s function into the formula (D.1), the calculation of the topological invariant would reduce to the weakly coupled calculation using $G^{-1}(0,k)$ as an effective Hamiltonian. In the process to show that tuning $\lambda$ from zero to one is a smooth process, they have used the fact that $G(i\omega,k,\lambda)$ does not have zero eigenvalues at nonzero $\omega$, seemingly introducing a constraint that the Green’s function has to have no poles in the imaginary frequency space other than $\omega=0$.
Let us then move to the calculation of the topological invariant for a semimetal. In our case, to calculate the topological invariant for a semimetal, which is a gapless topological state, the topological invariant should be always calculated in a surface/circle surrounding the node, where the node is a zero point of the Green’s functions. Thus there is a small modification to that definition of the universal definition of the topological invariant, which now becomes
$N=\frac{1}{24\pi^{2}}\int d\omega\oint d^{2}k\text{Tr}[\epsilon^{\mu\nu\rho}G\partial_{\mu}G^{-1}G\partial_{\nu}G^{-1}G\partial_{\rho}G^{-1}].$ (D.3)
The integral on $k$ is restricted to the small surface enclosing the Weyl node in the momentum space while the integral in $\omega$ covers the whole real $\omega$ axis. This integration in the $k$ space automatically avoids the singular points or poles of the Green’s functions as the enclosing surface could be chosen to be extremely close to the Weyl point.
Note that at the same time, as we have explained, in the original work for topological Hamiltonian, the authors introduced imaginary frequency Green’s functions for their calculational convenience. In our case, it should be more straightforward and more justified to directly use the real frequency Green’s function. The reasons are that: the real-frequency Green’s function directly probes the physical spectral function, making it ideal for studying the low-energy physics of Weyl points and their topological properties; while the imaginary-frequency Green’s function is less directly tied to physical observables, as it represents the system in a Euclidean framework. Analytic continuation to the real axis may be needed to extract physical properties, which introduces numerical errors.
For strongly interacting Weyl semimetals, imaginary frequency Green’s functions would be much easier to calculate though not so relevant for physical systems, however, holography solves this problem so that we could use directly the real frequency Green’s functions. Therefore, in the formula for the calculation of topological invariants for holographic semimetals, the frequency should be only integrated in the real space. Thus as long as there are no poles/singlar points in real frequency space at $\omega\neq 0$, which is indeed true, the topological Hamiltonian approach is justified using the tuning $\lambda$ method. Especially, the integration in the $k$ space is performed over an extremely small enclosing surface around the Weyl point. This ensures that it is quite easy to verify that there are no singular points or poles for the Green’s functions within the entire frequency range when k is restricted to be extremely close to the Weyl point.
Thus, we could instead define the following real frequency version of the smooth function
$G(w,k,\lambda)=(1-\lambda)G(w,k)+\lambda[\omega+G^{-1}(0,k)]^{-1}.$ (D.4)
As long as $G(w,k,\lambda)$ is regular for nonzero $\omega$ in the real frequency space, the topological invariant calculated for $\lambda=1$ would be the same for $\lambda=0$, i.e. we could use the topological Hamiltonian method to use $G^{-1}(0,k)$ as the effective Hamiltonian for topological invariants.
Therefore, in our work, we do not require the Green’s function to have no poles in the imaginary frequency space. For our use, the topological Hamiltonian method is justified and the topological invariants are robust and well-captured by the zero-frequency limit of the Green’s function. This also reflects the fact that topology is encoded in the global properties of the Green’s function, and the zero-frequency slice often retains this essential topological information.